A rectangular gasoline tank can hold 63 kg of gasoline when full. What is the depth of the tank if it is 0.335-m wide by 0.749-m long? (Assume that


Round your answer to 3 decimal places


1
Expert's answer
2021-04-16T07:26:49-0400

Find volume:


"V=abc."

On the other hand, it is


"V=\frac m\rho."

Find depth "c":


"abc=\frac m\rho,\\\space\\
c=\frac{m}{\rho ab}=\frac{63}{800·0.335·0.749}=0.314\text{ m}."